OpenStudy (anonymous):

evaluate. log(1/square root 1000)

OpenStudy (lgbasallote):

no base...so it means it's a common log where the base is automatically 10 since we do not know the answer, we'll equate it to x.so.. log_10 (1/sqrt 1000) = x what do you think happens next?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

how so?

OpenStudy (anonymous):

root 1000 is 100 \[\log_{10} 1/\sqrt{10000}=-\log_{10} 100=-2\]

OpenStudy (anonymous):

umm.. square root 10000 is 100. my problem is square root 1000. thank you tho!

OpenStudy (anonymous):

oops wait...:(

OpenStudy (anonymous):

can it be -log10 square root 10?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

ans -3/2

OpenStudy (anonymous):

wow... -3/2?

OpenStudy (lgbasallote):

log_10 (1/sqrt 1000) = x 10^x = 1000^-1/2 1000 = 10^3 so... 10^x = (10^3)^-1/2 10^x = 10^-3/2 x = -3/2 according to the same base thing (if a^m = a^n then m=n)